# CloudHeatMap: Heatmap-Based Monitoring for Large-Scale Cloud Systems

### v1

@@ Abstract
Cloud computing is essential for modern enterprises, requiring robust tools to monitor and manage Large-Scale Cloud Systems (LCS). Traditional monitoring tools often miss critical insights due to the complexity and volume of LCS telemetry data. This paper presents CloudHeatMap, a novel heatmap-based visualization tool for near-real-time monitoring of LCS health. It offers intuitive visualizations of key metrics such as call volumes, response times, and HTTP response codes, enabling operators to quickly identify performance issues. A case study on the IBM Cloud Console demonstrates the tool’s effectiveness in enhancing operational monitoring and decision-making. A demonstration is available at https://www.youtube.com/watch?v=3u5K1qp51EA.
Index Terms:
Cloud computing, Heatmaps, Cloud monitoring
I Introduction
@@
Cloud computing has become a critical infrastructure for businesses, providing scalability and cost reduction. However, as Cloud adoption grows, monitoring these systems becomes increasingly complex [1, 2, 3]. Large-Scale Cloud Systems (LCS) comprise numerous interconnected microservices distributed across data centers (see [4] for an example), generating vast amounts of telemetry data [5] essential for understanding system health and performance. The sheer volume and complexity of these data challenge traditional monitoring tools, which often fail to provide timely, actionable insights [1, 2, 3]. This lack can lead to undetected performance issues, inefficient resource allocation, and, ultimately, system failures that impact service delivery [1, 2, 3].
@@
Traditional Cloud monitoring tools (e.g., DataDog, Dynatrace) offer limited root cause visibility and are often reactive, alerting after issues impact the system. In dynamic Cloud environments, delayed responses can cause significant downtime and resource inefficiency. According to Google’s Site Reliability Engineering (SRE) principles, the four Golden Rules of Monitoring — traffic, errors, latency, and saturation — are essential metrics for assessing LCS health [6]. However, most monitoring tools focus on the interface-level analysis of system components [7], detecting issues but not providing deeper insights or predicting problems. While some advanced monitoring systems offer visual dashboards [8], they often lack support for exploratory health analysis. Operators need to delve into data patterns and understand component behaviours under varying conditions [9]. In today’s dynamic Cloud environments, early and proactive detection are critical [10, 4, 11, 12, 13, 14]; reactive approaches are insufficient. Thus, advanced monitoring solutions are needed that provide real-time insights and enable exploratory analysis to help Cloud Operators (Ops) teams identify issues before they escalate [1, 2, 3].
@@
The authors, drawing from their experience designing and using AIOps tools [1, 2, 3, 10, 4], can attest that while these tools are useful, they have limitations in root cause analysis and detecting persistent issues. Long-standing abnormalities may escape automated detection, requiring human operators to assess the broader system and make strategic decisions beyond local failures.
@@
We present CloudHeatMap [^1] The tool is available via https://github.com/sohanasarah/CloudHeatMap and includes $\approx 24$ hours of the anonymized real data., a novel heatmap-based [^2] A grid heatmap maps data points as colours in two dimensions, where colour intensity represents data magnitude [15], helping identify patterns and relationships in large datasets, allowing for quicker decision-making. visualization tool leveraging telemetry data from microservices to visualize key performance metrics like call volumes, response times, and HTTP (as well as custom) response codes.
@@
We drew inspiration from the classical annunciator panels [16], that were commonly used in power plants and aircraft control boards. These panels feature a grid of lights or buttons, each representing a system parameter. When a parameter changes, the light or button updates its color or state, enabling operators to efficiently monitor processes and respond to alerts.
@@
As we empirically found, heatmaps are valuable because they condense large amounts of data from complex systems into a simple grid, where color highlights anomalies, allowing operators to quickly and intuitively detect unusual behaviour.
@@
CloudHeatMap offers a near-real-time system health overview, enabling operators to quickly identify hot spots and potential issues. The tool helps Cloud Ops make informed decisions about resource allocation, configuration changes, and problem resolution, ultimately enhancing Cloud services’ reliability and efficiency. It also supports temporal analysis, helping operators track system behaviour over time.
@@
To evaluate tools’s effectiveness, we conducted a case study on the IBM Cloud Console, an LCS with significant operational demands. Results show that CloudHeatMap  enhances Cloud Ops’ ability to monitor and maintain system health, providing insights previously difficult to obtain with traditional methods. This research contributes to Cloud computing by offering a scalable and practical solution for monitoring LCS.
II Methodology
This research methodology [^3] For further details, please refer to the M.Sc. thesis by the first author [17]. involves the design, development, and evaluation of CloudHeatMap  tailored for monitoring the health of an LCS. The process is divided into key phases: system analysis, data collection, data processing, and visualization design. The diagram in Fig. 1 shows the sequence of actions required from data collection to heatmap generation.
Note that CloudHeatMap focuses on visualization (Section II-D) and operates independently of data collection and processing (Sections II-B and II-C). These phases are included as part of a representative example to demonstrate a complete data pipeline for context. Consequently, CloudHeatMap can be used with any data source, as long as the data is provided in the expected format.
Figure 1: System diagram.
II-A System Analysis
@@
The first step involved analyzing the LCS under study, the IBM Cloud Console, a critical component of IBM’s Cloud infrastructure [4], serving as the web front-end and orchestrator for the IBM Cloud platform [18]. Currently, the system consists of around 125 microservices deployed worldwide. Each microservice generates millions of log records daily, encompassing various operational metrics, including HTTP response codes, call volumes, and response times [4].
@@
II-B Data Collection
The data collection phase focused on capturing telemetry data from the IBM Cloud Console. In this context, telemetry data refers to the metrics, logs, and traces emitted by the IBM Cloud Console microservices. The traces contain detailed information about each request processed by the microservices, including timestamps, response codes, response times, and the identities of caller and callee microservices [^4] A user request triggers calls between microservices. The initiating microservice is the caller, and the receiving one is the callee..
We designed and implemented a robust data collection and analysis pipeline, leveraging our previous work [19, 20, 4]. The pipeline was designed to be data-agnostic and could persistently store data in IBM COS for near-real-time analysis. A Pub/Sub [21] mechanism managed the high data volume, with each microservice publishing telemetry to a Redis Pub/Sub pipeline. Microservices running as containers, managed by Kubernetes, extracted, transformed, and loaded (ETL) the necessary data into persistent storage system.
A sub-pipe connected to the Redis Pub/Sub captured traces in the Zipkin format [22] (with support for extending to formats such as OpenTrace [23] and OpenTelemetry [24]). Telemetry was collected at 1-minute intervals to provide near-real-time information. Due to the large data volume, we filtered the data during collection based on recommendations from the IBM Console team.
The sub-pipe includes an ETL module, extracting essential fields such as timestamps, trace IDs, microservice names, HTTP status codes, and response times from the raw telemetry data. The extracted data are then transformed into arrays capturing computed statistics for each microservice across all production data centers and for caller-callee microservice pairs. These statistics included metrics such as call volumes, average response times, and the distribution of HTTP response codes (see Table I for the complete list).
TABLE I: Metrics generated from the extracted data.
@@ II-B Data Collection
HTTP Status Code Percentage and value of HTTP response codes
Finally, the transformed data were stored as a nested JSON object with indexed timestamps, facilitating efficient time series analysis. The structured data were then used to generate heatmaps providing an intuitive visualization of system health metrics.
II-C Data Processing
Once collected, the raw data needed processing and aggregation into defined intervals suitable for visualization. The data processing phase involved several steps:
Data Aggregation
We combined each 1-minute interval of data stored in IBM Cloud Object Storage (COS) [25] over specific periods (e.g., 24 hours) using IBM Cloud Functions [^5] Replaced by IBM Cloud Code Engine [26] at the time of writing. [27]. A custom action aggregated new files into a combined file with a maximum size of 1 MB per file. After aggregation, functions such as sum, average, maximum, minimum, and standard deviation generated summary statistics for each microservice.
Data Transformation
Aggregated data were transformed into a format suitable for heatmap visualization, converting the aggregated metrics into two-dimensional arrays where rows represented data centers or caller microservices, and columns represented monitored microservices. The transformation included resampling data to desired time intervals (e.g., 15 minutes, hourly) using aggregation functions (e.g., compound mean and standard deviation, see [17] for details).
Data Storage
The processed data were stored in IBM COS, which provided a scalable and secure environment for managing the large datasets the IBM Cloud Console generated. The data retrieval module, written in Python, is connected to the IBM COS API to read objects from the storage bucket within a user-defined time window. The retrieved JSON data was converted into a two-dimensional DataFrame for further processing.
II-D Visualization Design
@@
The core part of the tool involved designing and implementing the heatmap visualization tool using the Plotly Dash library [28, 29].
@@
Figure 2: A heatmap comparing data centers and microservices, useful for assessing whether an issue has a greater impact on data centers.
@@
Figure 3: A heatmap showing caller-callee microservices, useful for identifying issues in the interactions between microservices.
@@
CloudHeatMap visualizes call volumes, response times, and HTTP response codes across the IBM Cloud Console’s microservices and data centers. Figs. 2 and 3 show sample heat maps using actual anonymised data (included with the demo code). Fig. 2 illustrates the relationship between data centers and services, with microservice names on the $x$-axis and data centers on the $y$-axis. Fig. 3 depicts the caller-callee interactions, with callee microservices on the $x$-axis and caller microservices on the $y$-axis. Hovering with a mouse over a specific tile in the heatmap displays a window showing the numeric value of the tile along with its corresponding $x$ and $y$ coordinates, enhancing clarity. Key features include:
Interactive User Interface
@@
The tool offers a highly interactive UI with multiple filters, allowing customization based on graph type (data centers-services or caller-callee pairs), metrics (e.g., call volumes or average response time), and HTTP status codes. Users can switch between viewing absolute values or percentages and filter by value type and range.
Temporal Analysis
@@
Users can play an animation displaying evolution of metrics over time, crucial for identifying trends and episodic issues that static views might miss. The first frame of the animation presents an aggregated view of the selected time period, offering a comprehensive snapshot, with subsequent frames showing behavioral changes over time.
@@
CloudHeatMap  is designed to highlight “hot spots” where likely, indicated by areas of higher colour intensity on the heatmap (e.g., indicating higher call volumes, longer response times, or higher error rates). This helps operators quickly identify and address potential issues.
III Evaluation
@@
We assessed CloudHeatMap through a two-month case study on IBM Cloud Console (discussed in Section II-A) in collaboration with the IBM Cloud Ops team, focusing on its ability to monitor and improve IBM Cloud Console’s performance. For details of the evaluation of use cases, see [17].
III-A Rate Limiting Detection
@@
Rate limiting is a concern in Cloud environments, where microservices must operate within specified limits to prevent service degradation. CloudHeatMap analyzed HTTP 429 “Too Many Requests” response codes across microservices revealing that one microservice consistently exceeded acceptable rate limits across data centers, with peak values reaching 40% in some time intervals. This visualization allowed the Ops team to identify and address the issue promptly.
III-B Detecting Components’ Errors
@@
III-C Detecting Performance Degradation
User satisfaction is closely tied to interface responsiveness, with response times over 2 seconds degrading experience [30]. CloudHeatMap  identified (using caller-callee view) microservices contributing to latency issues, such as “Dashboard-broker” and “Preferences” interacting with a “Cloudant” database instance. Early detection allowed the Ops team to intervene before issues escalated into a critical failure.
III-D Re-architecting Hot-spots
@@ III-E Cost Savings
IV Discussion
CloudHeatMap  equipped the IBM Ops team with actionable insights that traditional monitoring tools could not, revealing hidden issues and distinguishing between persistent and episodic problems. This proactive monitoring approach provides a more comprehensive view of system health, unlike the retrospective analyses of other platforms like Datadog and Dynatrace. Additionally, CloudHeatMap bridges the gaps left by machine-learning-based tools, which, despite their sophistication, often struggle with exploratory analysis [31].
These use cases align with some of Musa’s Operational Profiles analysis [32], where observing client usage helps identify problematic components for testing and re-architecting, especially those frequently accessed or with high response times.

### v2

@@ Abstract.
Cloud computing is essential for modern enterprises, requiring robust tools to monitor and manage Large-Scale Cloud Systems (LCS). Traditional monitoring tools often miss critical insights due to the complexity and volume of LCS telemetry data. This paper presents CloudHeatMap, a novel heatmap-based visualization tool for near-real-time monitoring of LCS health. It offers intuitive visualizations of key metrics, such as call volumes and response times, enabling operators to quickly identify performance issues. A case study on the IBM Cloud Console demonstrates the tool’s effectiveness in enhancing operational monitoring and decision-making.
1. Introduction
@@
Cloud computing has become a critical infrastructure for businesses, providing scalability and cost reduction. However, as Cloud adoption grows, monitoring these systems becomes increasingly complex (Pourmajidi et al., 2017, 2019, 2021). Large-Scale Cloud Systems (LCS) comprise numerous interconnected microservices distributed across data centers (see (Islam et al., 2021, 2024) for an example), generating vast amounts of telemetry data (Miranskyy et al., 2016) essential for understanding system health and performance. The sheer volume and complexity of these data challenge traditional monitoring tools, which often fail to provide timely, actionable insights (Pourmajidi et al., 2017, 2019, 2021). This lack can lead to undetected performance issues, inefficient resource allocation, and, ultimately, system failures that impact service delivery (Pourmajidi et al., 2017, 2019, 2021).
@@
Traditional Cloud monitoring tools (e.g., DataDog, Dynatrace) offer limited root cause visibility and are often reactive, alerting after issues impact the system. In dynamic Cloud environments, delayed responses can cause significant downtime and resource inefficiency. According to Google’s Site Reliability Engineering (SRE) principles, the four Golden Rules of Monitoring — traffic, errors, latency, and saturation — are essential metrics for assessing LCS health (Ewaschuk, 2016). However, most monitoring tools focus on the interface-level analysis of system components (Schoonjans et al., 2015), detecting issues but not providing deeper insights or predicting problems. While some advanced monitoring systems offer visual dashboards (Cinque et al., 2022), they often lack support for exploratory health analysis. Operators need to delve into data patterns and understand component behaviours under varying conditions (Baciu et al., 2017). In today’s dynamic Cloud environments, early and proactive detection are critical (Islam and Miranskyy, 2020; Islam et al., 2021; Hrusto et al., 2022, 2023, 2024; Zhong et al., 2024); reactive approaches are insufficient. Thus, advanced monitoring solutions are needed that provide real-time insights and enable exploratory analysis to help Cloud Operators (Ops) teams identify issues before they escalate (Pourmajidi et al., 2017, 2019, 2021).
@@
The authors, drawing from their experience designing and using AIOps tools (Pourmajidi et al., 2017, 2019, 2021; Islam and Miranskyy, 2020; Islam et al., 2021, 2024), can attest that while these tools are useful, they have limitations in root cause analysis and detection of persistent issues. Long-standing abnormalities can escape automated detection, requiring human operators to assess the broader system and make strategic decisions beyond local failures.
@@
We present CloudHeatMap [^1] The tool is available via (Sohana et al., 2025) and includes $\approx 24$ hours of synthetic data. A demonstration is available at (Sohana, 2025)., a novel heatmap-based [^2] A grid heatmap maps data points as colours in two dimensions, where colour intensity represents data magnitude (Wilkinson and Friendly, 2009), helping identify patterns and relationships in large datasets, allowing for quicker decision-making. visualization tool leveraging microservices telemetry data to visualize key performance metrics, such as call volumes and response times, with filtering options for HTTP and custom response codes.
@@
We drew inspiration from the classical annunciator panels (isa, 1979), that were commonly used in power plants and aircraft control boards. These panels feature a grid of lights or buttons, each representing a system parameter. When a parameter changes, the light or button updates its colour or state, enabling operators to monitor processes efficiently and respond to alerts.
@@
As we empirically found, heatmaps are valuable because they condense large amounts of data from complex systems into a simple grid, where colour highlights anomalies, allowing operators to quickly and intuitively detect unusual behaviour.
@@
CloudHeatMap (described in Section 2) offers a near-real-time system health overview, enabling operators to quickly identify hot spots and potential issues. The tool helps Cloud Ops make informed decisions about resource allocation, configuration changes, and problem resolution, ultimately enhancing Cloud services’ reliability and efficiency. It also supports temporal analysis, helping operators track system behaviour over time.
@@
To evaluate tools’s effectiveness, we conducted a case study (given in Section 3) on the IBM Cloud Console, an LCS with significant operational demands. Results show that CloudHeatMap  enhances Cloud Ops’ ability to monitor and maintain system health, providing insights previously difficult to obtain with traditional methods. This research contributes to Cloud computing by offering a scalable and practical solution for monitoring LCS.
2. Tool Description
CloudHeatMap user interface, described in Section 2.2, operates independently of the data collection method. Thus, for different application deployments, the telemetry collection process may vary. However, as long as the data are converted into the specified format — outlined in Section 2.3 — the tool can visualize the data effectively. Figure 1 illustrates the sequence of actions involved, from data collection to heatmap generation.
For completeness, Section 2.1 provides a brief overview of the data collection and processing mechanisms used in our studies. Note that the complexity of the collection process can vary based on the requirements and preferences of the organization responsible for maintaining and operating a given software.
Figure 1. System diagram.
2.1. Original System and Data Harvesting
2.1.1. System Analysis
@@
The first step involved analyzing the LCS under study, the IBM Cloud Console, a critical component of IBM’s Cloud infrastructure (Islam et al., 2021), serving as the web front-end and orchestrator for the IBM Cloud platform (ibm, [n. d.]). Currently, the system consists of around 125 microservices deployed worldwide. Each microservice generates millions of log records daily, encompassing various operational metrics, including response codes, call volumes, and response times (Islam et al., 2021).
@@
2.1.2. Data Collection
Our data collection focuses on telemetry from the IBM Cloud Console, comprising metrics, logs, and traces emitted by microservices. Traces include essential request details such as timestamps, response codes, response times, and caller-callee microservice interactions [^3] A user request triggers calls between microservices. The initiating microservice is the caller, and the receiving one is the callee.. A robust, data-agnostic pipeline was designed using Redis Pub/Sub (Birman and Joseph, 1987) for handling high data volumes and IBM Cloud Object Storage (COS) (COS, [n. d.]) for persistence.
Telemetry data, published at 1-minute intervals, was captured in Zipkin format (zip, [n. d.]) (with support for extending to formats such as OpenTrace (ope, [n. d.]) and OpenTelemetry (Ope, [n. d.])). An extract, transform, and load (ETL) module extracted key fields (e.g., trace IDs, HTTP status codes) and generated summary statistics for the metrics in Table 1.
The processed data was stored as nested JSON with indexed timestamps (see Section 2.3 for format details) for efficient time series analysis. This structure enabled the creation of system health visualizations, including heatmaps.
Table 1. Metrics generated from the extracted data.
@@ 2.1.2. Data Collection
Status/Response Code Percentage and value of status/response codes
2.1.3. Data Processing
Collected data was aggregated into user-defined intervals (e.g., 24 hours) via IBM Cloud Functions [^4] Replaced by IBM Cloud Code Engine (Cod, [n. d.]) at the time of writing. (fun, [n. d.]), producing files capped at 1 MB. Summary statistics, including sums and averages, were computed for each microservice. The aggregated data were resampled at desired time intervals (e.g., 15 minutes, hourly) using aggregation functions (e.g., compound mean and standard deviation, see (Sohana, 2022) for details).
Finally, the processed data were stored in IBM COS, providing secure, scalable access. A Python-based retrieval module interfaced with the COS API to load JSON data within specified time windows, converting them into two-dimensional matrices for further analysis.
2.2. Visualization Design
@@
The core part of the tool involved designing and implementing the heatmap visualization tool using the Plotly Dash library (das, [n. d.]; plo, [n. d.]).
@@
Figure 2. A heatmap comparing data centers and microservices, useful for assessing whether an issue has a greater impact on data centers.
@@
Figure 3. A heatmap showing caller-callee microservices, useful for identifying issues in the interactions between microservices.
@@
CloudHeatMap visualizes call volumes and response times across the IBM Cloud Console’s microservices and data centers. Figures 2 and 3 show sample heat maps using actual anonymised data. Figure 2 illustrates the relationship between data centers and services, with microservice names on the $x$-axis and data centers on the $y$-axis. Figure 3 depicts the caller-callee interactions, with callee microservices on the $x$-axis and caller microservices on the $y$-axis. Hovering with a mouse over a specific tile in the heatmap displays a window showing the numeric value of the tile along with its corresponding $x$ and $y$ coordinates, enhancing clarity. Two key features are as follows.
  1. (1)
@@
    Interactive User Interface. The tool offers a highly interactive UI with multiple filters, allowing customization based on graph type (data centers-services or caller-callee pairs), metrics (e.g., call volumes or average response time), and return codes. Users can switch between viewing absolute values or percentages and filter by value type and range.
  2. (2)
@@
    Temporal Analysis. Users can play an animation displaying evolution of metrics over time, crucial for identifying trends and episodic issues that static views might miss. The first frame of the animation presents an aggregated view of the selected time period, offering a comprehensive snapshot, with subsequent frames showing behavioral changes.
@@
CloudHeatMap  is designed to highlight likely “hot spots” with areas of higher colour intensity on the heatmap (e.g., indicating higher call volumes, longer response times, or higher error rates). This helps operators quickly identify and address potential issues.
2.3. Data formatting for visualization
CloudHeatMap consumes the data in the format shown in Listing 1. For each timestamp (formatted as Unix timestamp in milliseconds), two dictionaries are recorded: datacenter_services and caller_callee_pairs, corresponding to the data visualized in Figures 2 and 3, respectively.
2.3.1. Data for Figure 2
The datacenter_services structure contains two primary identifiers: app_instance_id (mapped to data center in our case) and microservice_id.
The microservice_id corresponds to individual microservices within the application. In Figure 2, the $y$-axis represents the app_instance_id, while the $x$-axis represents the microservice_id.
For each app_instance_id and microservice_id pair, the data captures return codes. For each return code, we record a dictionary of statistics represented as pairs of the statistics name (e.g., stats_name_1 or min) and a numeric value. These pairs are visualized as a heatmap in the figure.
2.3.2. Data for Figure 3
The caller_callee_pairs structure is similar to datacenter_services, but it captures relationships between pairs of microservices. Specifically, it records the interactions between a caller microservice (caller_microservice_id) and a callee microservice (callee_microservice_id), rendered on $y$-axis and $x$-axis of Figure 3, respetively.
For each caller-callee pair, the data captures return codes and their associated statistics, which are then rendered as a heatmap to visualize the interactions.
Listing 1: JSON data structure injested by CloudHeatMap.
1{"timestamp": [
2 {
3 "datacenter_services": {
4 "app_instance_id": {
5 "microservice_id": {
6 "return_code_1": {
7 "stats_name_1": 1.0, "stats_name_2": 0.1},
8 "return_code_2": {
9 "stats_name_1": 5.0, "stats_name_2": 6.2}
10 }}},
11 "caller_callee_pairs": {
12 "caller_microservice_id_1": {
13 "callee_microservice_id_1": {
14 "return_code_1": {
15 "stats_name_1": 1.1, "stats_name_2": 0.1}},
16 "callee_microservice_id_2": {
17 "return_code_1": {
18 "stats_name_1": 1.5, "stats_name_2": 0.5}}
19 }}}]}
2.3.3. Generalization Notes
The tool supports rendering telemetry from any application. The app_instance_id identifies an application instance. The physical location of the application can vary based on the user’s deployment strategy; however, CloudHeatMap remains agnostic to these deployment specifics. CloudHeatMap can monitor single or multiple instances of the same application, each distinguished by a unique app_instance_id.
Return codes can be any string-convertible format, not limited to HTTP status codes. If the application does not provide return codes, a placeholder return code can be hardcoded.
Statistical metrics can vary according to user preferences, provided that the statistic’s name can be represented as a string and its value as a floating-point number.
The JSON files can be stored either in local storage (e.g., as demonstrated in our demo) or in cloud storage (e.g., as implemented in production)—depending on the user’s preference.
3. Evaluation
@@
We assessed CloudHeatMap through a two-month case study on IBM Cloud Console (discussed in Section 2.1.1) in collaboration with the IBM Cloud Ops team, focusing on its ability to monitor and improve IBM Cloud Console’s performance. For details of the evaluation of use cases, see (Sohana, 2022).
3.1. Rate Limiting Detection
@@
Cloud microservices must remain within the rate limits to avoid service degradation. CloudHeatMap revealed one microservice frequently exceeding its rate limit, with HTTP 429 “Too Many Requests” responses peaking at 40% across data centers in some time intervals. This visualization helped the Ops team quickly identify and address the issue.
3.2. Detecting Components’ Errors
@@
3.3. Detecting Performance Degradation
User satisfaction is closely tied to interface responsiveness, with response times over two seconds degrading experience (Anderson, 2022). CloudHeatMap  identified (using caller-callee view) microservices contributing to latency issues, such as “Dashboard-broker” and “Preferences” interacting with a “Cloudant” database instance. Early detection allowed the Ops team to intervene before issues escalated into a critical failure.
3.4. Re-architecting Hot-spots
@@ 3.5. Cost Savings
4. Discussion
CloudHeatMap is a simple yet effective tool that equipped the IBM Ops team with actionable insights that traditional monitoring tools could not, revealing hidden issues and distinguishing between persistent and episodic problems. This proactive monitoring approach provides a more comprehensive view of system health, unlike the retrospective analyses of other platforms like Datadog and Dynatrace. Additionally, CloudHeatMap bridges the gaps left by machine-learning-based tools, which, despite their sophistication, often struggle with exploratory analysis (Milo and Somech, 2020). By allowing operators and designers to reveal complex patterns requiring intervention, it enhances system comprehension and situational awareness.
These use cases align with some of Musa’s Operational Profiles analysis (Musa, 1993), where observing client usage helps identify problematic components for testing and re-architecting, especially those frequently accessed or with high response times.